OpenStudy (monteskellie):

Solve x + 3y = 9 3x − 3y = −13

OpenStudy (unklerhaukus):

Add the left hand sides together, and also add the right hand sides together ( x + 3y ) + [ 3x - 3y ] = ( 9 ) + [ -13 ]

OpenStudy (unklerhaukus):

Solve this equation for x

OpenStudy (unklerhaukus):

Plug this x you find into either of the original equations and solve for y.

OpenStudy (wolf1728):

a) x + 3y = 9 b)3x − 3y = −13 Add the equations 4x = -4 x = -1 Substitute the value of x into equation "a" a) -1 +3y = 9 Check your answers with this calculator: http://www.1728.org/unknwn2.htm